# Modeling the formation of N2 and CH4 frost on Pluto’s slopes

### v1

@@ 2.3.2 Modifications of the parameterization for the Pluto Planetary Climate Model
    The visible fluxes should be computed considering the effects of aerosols/hazes in the atmosphere. To do so, one can compute this solar irradiance on a flat surface, as traditionally done by radiative transfer algorithms (e.g., Forget et al., 2017; Bertrand & Forget, 2017) and then derive it for any slope based on geometric considerations (see for instance Eqs. 3 and 4 in Spiga & Forget, 2008). One should also estimate the scattered irradiance on slopes by the aerosols. However, this requires constructing a complete parameterization using radiative transfer models and is out of the scope of this paper and left as future work.
@@
    The radiative effect of the atmosphere at infrared wavelengths should also be included in the surface energy budget. To do so, one can derive the downwelling infrared radiance $F_{\rm{rad,lw,flat}}$ for a flat surface using radiative transfer models (e.g., Forget et al., 2017; Bertrand & Forget, 2017), and then add the term $\sigma_{s}F_{\rm{rad,lw,flat}}$ in Eq. 5.
@@
    When N2 is condensing, Eq. 6 should be replaced by (Forget et al., 2017):
    $\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)=\frac{\rho c_{s}}{\Delta t}\frac{\left(T_{\rm{cond}}-T_{0}\right)}{L_{\rm{N}_{2}}+c_{p}\left(T_{1}-T_{\rm{cond}}\right)}$ (11)
@@ 2.3.3 Validation and limits of the model
Validating the parameterization on Pluto is challenging due to the paucity of measurements, especially at the slope scale, and the strong sensitivity of ice dynamics to surface properties as highlighted by previous studies (e.g., Bertrand et al., 2018, 2019). Given the large validation of the parameterization on Mars using various datasets performed by Lange et al. (2023), we will consider that the application of the parameterization on Pluto is reliable. Yet, our parameterization has several limits that we acknowledge here. First, as mentioned previously, the microclimate of East-West-facing slopes is not modeled here. Second, shadows cast by surrounding terrain (as opposed to slope-induced self-shadowing) are neglected. This omission prevents the representation of cold traps, e.g., at the bottom of a ridge/canyon. Modeling such shadowing would require ray-tracing computation for each point modeled here at each time of the year, leading to a too significant increase in the computation time. Third, slope winds are not explicitly modeled: wind speeds on sub-grid slopes are assumed to follow large-scale winds simulated in the Pluto GCM or prescribed in the VTM. However, katabatic winds might induce stronger winds on crater rims, as for instance modeled on the flanks of the Sputnik Planitia basin (Forget et al., 2017; Bertrand et al., 2020). These slope winds should impact the sublimation rate of methane deposits as shown by Eq. 7. In the absence of slope-wind parameterization in our model, we test the sensitivity of our results to these slope winds by varying $U$ in Eq. 7 (see Table 1). Finally, a shared atmospheric column is assumed for all sub-grid slopes, implying effective mixing by large-scale or regional circulation. Testing this would require mesoscale or LES modeling, beyond the scope of this study.
2.4 Description of the simulations performed
@@
As mentioned in Section 2.2, several parameters must be adjusted to reproduce the surface pressure and CH4 volume mixing ratio observed by stellar occultations and high-resolution spectroscopy (e.g., Meza et al., 2019; Lellouch et al., 2009). The default parameters listed in Table 1 provide a good representation of the seasonal evolution of surface pressure and CH4 volume mixing ratio, in agreement with the observations of Meza et al. (2019) and Lellouch et al. (2015) (see, for instance, Figure 2 of Bertrand & Forget, 2016). Given the sensitivity of the model to the inputs listed in Table 1, we investigate the robustness of our results by varying the input parameters. However, since modifying the albedo or emissivity of Pluto’s main ice reservoirs would affect global surface pressure and CH4 levels, influencing frost condensation/sublimation on slopes, we chose to vary the surface and ice properties in our sensitivity only within the longitudes of the Cthulhu region, between latitudes 40°N and 40°S. The baseline values were retained for the rest of the surface to maintain realistic pressure and CH4 cycles.
@@
Table 1: Input parameters for the Pluto VTM and the values adopted in our sensitivity study. In the latter case, parameter values are varied only within the Cthulhu region, while elsewhere they remain fixed at their baseline values. The transition between the albedo of the bare surface and that of a CH4-frosted surface follows the parameterization of Bertrand et al. (2020). The values listed in this table correspond to the maximum albedo that a frosted surface can reach, typically for deposits thicker than $\sim$ 1 mm. TI refers to the thermal inertia, the subscript d corresponds to the diurnal thermal inertia, i.e., first centimeters of the surface, and the subscript s refers to the seasaonal thermal inertia (depths deeper than $\sim$ 5 cm.
For each set of parameters, the model was run for 100 Pluto’years to equilibrate the subsurface temperatures on each sub-grid slope, with slope condensation/sublimation disabled to avoid the formation or disappearance of thick glaciers due to non-equilibrated temperatures. We then ran the model for an additional ten years, allowing the condensation of volatiles on each sub-grid slope. Only results from the final year are presented in this manuscript.
@@ 3.1 Frost mapping and statistics
Figure 4: CH4 integrated LEISA band depth maps from Schmitt et al. (2017) and Gabasova et al. (2021), overlaid with LORRI albedo map (Hofgartner et al., 2023). The white patches in the Cthulhu region, which are not colored, are considered to be CH4 frost (see the main text for explanation).
The geographic distribution of frost identified in this study, based on the methodology described in section 2.1 is presented in Figure 4. A comparison between the LEISA CH4 and N2 band depth maps at this location indicates that these frosts correspond to N2 ice deposits in Sputnik Planitia (northeastern part of Figure 4) and Elliot crater ($\sim$ 12°N, 139°E), and to CH4 frost elsewhere. CH4 frost is detected poleward of 10°N, but its identification at lower latitudes is more difficult, as the bright spots observed with LORRI are not always associated with a distinct signal in the CH4 integrated LEISA band depth. This bias arises from the resolution difference between the LORRI observations ($\sim$ 300 m per pixel, Weaver et al., 2008; Hofgartner et al., 2023) and the LEISA CH4 band depth maps ($\sim$ 2.7 km at best, Gabasova et al., 2021).
Nonetheless, several lines of evidence support the interpretation that these bright deposits are indeed CH4 frost. First, in the largest bright patches, a few pixels exhibit high CH4 integrated band depth values (up to 0.4). In such cases, the correlation between bright deposits and CH4 frost is robust, suggesting that the absence of a correlation elsewhere is primarily due to resolution differences between the instruments. Second, Earle et al. (2018b) identified CH4 ice deposits on slopes between 15°S and 15°N using MVIC data, which offers higher spatial resolution ($\sim$ 1 km Earle et al., 2022). Third, analysis of RGB images at these locations reveals that some bright, high-albedo slopes (Supplementary Figure 6 of Bertrand et al., 2020) are adjacent to similarly oriented, red, volatile-free slopes (Earle et al., 2022), indicating that the bright signatures are not solely due to illumination effects but also to the presence of surface ice. We therefore conclude that the bright deposits observed in Figure 4 within the Cthulhu region are CH4 frost. The LEISA phase index map (see Supplementary Figure 7b of Bertrand et al., 2020) suggests that these deposits are relatively pure CH4. Earle et al. (2022) further suggests that pure CH4 ice is observed on crater rims at low latitudes, while CH4 diluted with N2 may occur on the North-facing crater walls at lower elevations.
@@
Figure 5: Normalized count of A) the slope angle, B) the slope azimuth, C) the projected slope $\mu$ where CH4 frost is detected for all locations on Figure 4 (grey histogram) and only the Cthulhu region ($\pm$ 10° latitudes, blue histogram).
3.2 Simulation of frost formation and evolution on slopes
@@ 3.2.1 Baseline results
The seasonal evolution of CH4 and N2 frosts on flat surfaces, and on south- and North-facing slopes, using the baseline model parameters described in Table 1, is shown in Figure 6. These results highlight the significant influence of slope orientation on the condensation and sublimation of N2 and CH4. At high latitudes (Figures 6a, b), the amount of N2 frost forming on 30° North-facing slopes is about five times higher (three times for CH4) than on flat terrain. Furthermore, our model predicts perennial accumulation on 30°South-facing slopes, whereas only seasonal frost is predicted on flat surfaces. At 20°N and at the equator (Figures 6c–h), N2 and CH4 frosts form almost exclusively on slopes, with only negligible amounts on flat terrain. In the Southern hemisphere (10°S, Figures 6g, h), frost is again predicted on South-facing slopes and flat surfaces during Southern winter, but not during Northern winter, while our model also predicts frost formation on North-facing slopes.
@@
These patterns can be explained by the differential solar illumination received by surfaces with different orientations. As shown by Bertrand et al. (2020) (see their supplementary Figs. 8 and 9), 30° North-facing slopes in the Northern hemisphere receive about 35% less insolation than flat terrain during winter and spring (and conversely for 30° South-facing slopes). Volatiles should therefore accumulate preferentially on these cold traps relative to warmer surfaces. The seasonal evolution of frost deposits also follows the illumination cycle: frost forms in late autumn and winter, when slopes are in shadow, and sublimates in spring and summer, when they are exposed to sunlight. However, an important asymmetry is observed: frost never completely disappears from South-facing slopes, unlike North-facing slopes (although the latter is highly dependent on the assumed albedo, see section 3.2.2). In contrast, Bertrand et al. (2020) suggested, based on illumination computations, that frost on South-facing slopes should disappear after Southern spring, when these slopes warm. The main difference between our model and theirs is the inclusion of the seasonal CH4 cycle, which strongly affects the sublimation rate of CH4 frost. As shown by Bertrand & Forget (2016) (see their Fig. 2), the maximum atmospheric CH4 mixing ratio during Southern summer is five times higher than during Northern summer, because the atmosphere is then ten times thinner and provides less dilution of methane. As expressed in Eq. 7, the sublimation rate of CH4 frost is highly sensitive to the CH4 mixing ratio. This asymmetry results in a sublimation flux up to 3-6 times lower during the Southern summer than in the Northern summer, thereby favoring the formation of perennial CH4 frost deposits on South-facing slopes. There is also a seasonal asymmetry in the atmospheric -and methane- density $\rho$, with larger values during the Northern spring compared to the Southern spring. This induces stronger sublimation fluxes in the Northern spring on warm, North-facing slope, and, conversely, enhanced CH4 condensation and accumulation on cold, South=facing slopes at the same periods, again favoring the preservation of frost and the formation of a perennial deposit. Once such a perennial deposit forms, its high albedo reduces the absorbed energy and the annual mean surface temperature, which in turn promotes further frost accumulation, establishing a positive feedback. Finally, note that despite Pluto’s strong orbital eccentricity, no effect related to surface temperature is detected, underlining that the seasonal asymmetry in CH4 frost dynamics is mainly driven by variations in atmospheric pressure, density and methane mixing ratio rather than by thermal forcing. Indeed, as shown by Bertrand et al. (2020) (see their Figure S8), North and South-facing slopes receive nearly the same amount of solar insolation during spring and summer, when illuminated, and thus reach comparable temperatures. This results from Pluto’s equinox and perihelion occurring close in time, which produces polar nights of equal length and summers of similar duration and intensity in both hemispheres (Earle et al., 2017).
@@
Figure 6: Seasonal evolution of frost on 30° South-facing slopes (blue), flat surfaces (red), and 30° North-facing slopes (black) on Pluto, shown for latitudes 40°N (A, B), 20°N (C, D), 0°N (E, F), and 10°S (G, H). For each latitude, N2 frost thickness is shown in the left panels, and CH4 frost thickness in the right panels. The grey dashed line marks the date of the New Horizons flyby. For cases with perennial accumulations, thickness values have been rescaled for clarity. The frost thicknesses shown here are zonal averages of the model outputs between 40°E and 150°E, corresponding to the Cthulhu region.
@@ 3.2.2 Model sensitivity
Figure 7 illustrates the sensitivity of our model results to the various parameters described in Table 1. As expected, increasing the ice albedo enhances the stability of frost (either N2 or CH4) during spring and summer (by about 20 Earth years for the case of N2, Figure 7a, and $\sim$7 years for CH4, Figure 7b). The stability of CH4 frost is highly sensitive to its albedo (c.f. the albedo feedback effect described in Bertrand et al., 2020): increasing the ice albedo from 0.71 to 0.73 can lead to the formation of perennial ice deposits. Increasing the emissivity of the ice also enhances frost stability with a similar magnitude (Figures 7c and d), as it improves ice cooling and thus reduces the sublimation rate. Frost deposits are also sensitive to the thermal conductivity of the surface and underlying bedrock (here defined at depths greater than 5 cm), and thereby to thermal inertia. High thermal inertia surfaces are more conductive and release the heat stored during daytime/summer into the cold seasons, which reduces the amount of ice that can condense and hence decreases its stability. As highlighted by Eq. 7 and Figure 7g, the condensation/sublimation rate of CH4 ice is strongly controlled by wind speed: strong winds promote deposition but also accelerate sublimation, while weak winds inhibit both condensation and sublimation. However, in the latter case, given the thinness of the deposits, frost disappears faster than under high-wind conditions. Finally, slope angle also affects frost stability: steeper slopes receive lower mean annual insolation, leading to lower surface temperatures and thus favoring the formation of thicker and more persistent frost deposits (Figure 7h). Additional sensitivity tests (not shown) indicate that bare surface albedo and emissivity affect the amount of heat absorbed by the ground during the day and its cooling at night, respectively. These processes influence the diurnal formation of frost and potentially its seasonal persistence. The amount of CH4 in the atmosphere, and hence its transport, also affects the stability of frost deposits, as previously explained (see last paragraph).
@@
Our sensitivity experiments show that the timing of frost formation and disappearance is highly sensitive to the choice of key parameters, notably albedo, emissivity, diurnal thermal inertia, and regolith thermal inertia. These parameters have also been identified as critical for the evolution of volatile cycles across Pluto (e.g., Bertrand & Forget, 2016; Bertrand et al., 2018, 2019; Young et al., 2021). On Mars, the seasonality of CO2 and H2O frosts has been used to constrain the vertical stratification of the subsurface and its properties (notably the thermal inertia) (Vincendon et al., 2010a, b). Such an analysis is not possible here because of the strong multi-parameter sensitivity of our problem and the fact that we only have a single snapshot of frost dynamics, which leaves the system poorly constrained.
@@
Figure 7: Sensitivity of the N2 and CH4 frost thicknesses to the model inputs: A) N2 ice albedo; B) CH4 ice albedo; C) N2 ice emissivity; D) CH4 ice emissivity; E and F) thick lines: surface thermal inertia (first 5 cm), dashed lines: bedrock (below 5 cm) thermal inertia G) wind speed H) slope angles. All of the frost thickness are for a 30° facing North, located at 20°N.

### v2

@@ 2.3.2 Modifications of the parameterization for the planetary climate model for Pluto
    The visible fluxes should be computed considering the effects of aerosols/hazes in the atmosphere. To do so, one can compute this solar irradiance on a flat surface, as traditionally done by radiative transfer algorithms (e.g., Forget et al., 2017; Bertrand and Forget, 2017) and then derive it for any slope based on geometric considerations (see for instance Eqs. 3 and 4 in Spiga and Forget, 2008). One should also estimate the scattered irradiance on slopes by the aerosols. This requires constructing a complete parameterization using radiative transfer models and is beyond the scope of this paper. We leave this for future work.
@@
    The radiative effect of the atmosphere at infrared wavelengths should also be included in the surface energy budget. To do so, one can derive the downwelling infrared radiance $F_{\rm{rad,lw,flat}}$ for a flat surface using radiative transfer models (e.g., Forget et al., 2017; Bertrand and Forget, 2017), and then add the term $\sigma_{s}F_{\rm{rad,lw,flat}}$ in Eq. 5.
@@
    When N2 is condensing, Eq. 6 should be replaced by (Forget et al., 2017)
    $\frac{\partial m_{\rm{N}_{2}}}{\partial t}(t)=\frac{\rho c_{s}}{\Delta t}\frac{\left(T_{\rm{cond}}-T_{0}\right)}{L_{\rm{N}_{2}}+c_{p}\left(T_{1}-T_{\rm{cond}}\right)},$ (11)
@@ 2.3.3 Validation and limits of the model
Validating the parameterization on Pluto is challenging due to the paucity of measurements, especially at the slope scale, and the strong sensitivity of ice dynamics to surface properties as highlighted by previous studies (e.g., Bertrand et al., 2018, 2019). Given the extensive validation of the parameterization on Mars using various datasets performed by Lange et al. (2023), we will consider that the application of the parameterization on Pluto is reliable. Yet, our parameterization has several limits that we acknowledge here. First, as mentioned previously, the microclimate of east-west-facing slopes is not modeled here. Second, shadows cast by surrounding terrain (as opposed to slope-induced self-shadowing) are neglected. This omission prevents the representation of cold traps, for example, at the bottom of a ridge/canyon. Modeling such shadowing would require ray-tracing computation for each point modeled here at each time of the year, leading to a too significant increase in the computation time. Third, slope winds are not explicitly modeled: wind speeds on subgrid slopes are assumed to follow large-scale winds simulated in the Pluto GCM or prescribed in the VTM. Katabatic winds might induce stronger winds on crater rims, however, as was modeled on the flanks of the Sputnik Planitia basin (Forget et al., 2017; Bertrand et al., 2020), for instance. These slope winds should impact the sublimation rate of methane deposits as shown by Eq. 7. In the absence of slope-wind parameterization in our model, we test the sensitivity of our results to these slope winds by varying $U$ in Eq. 7 (see Table 1). Finally, a shared atmospheric column is assumed for all subgrid slopes, implying effective mixing by large-scale or regional circulation. Testing this would require mesoscale or LES modeling, beyond the scope of this study.
2.4 Description of the simulations
@@
As mentioned in Sect. 2.2, the optical properties of N2 and CH4 ices, and the thermal inertia of the substrate must be adjusted to reproduce the surface pressure and CH4 volume-mixing ratio observed by stellar occultations and high-resolution spectroscopy respectively (e.g., Hinson et al., 2017; Meza et al., 2019; Lellouch et al., 2015). The default parameters listed in Table 1 provide a good representation of the seasonal evolution of surface pressure and CH4 volume-mixing ratio, in agreement with the observations of Hinson et al. (2017), Meza et al. (2019), and Lellouch et al. (2015, see, for instance, Fig. 2 of ). Given the sensitivity of the model to the inputs listed in Table 1, we investigate the robustness of our results by varying the input parameters. Because modifying the albedo or emissivity of the main ice reservoirs on Pluto would affect the global surface pressure and CH4 levels, which in turn would affect the frost condensation/sublimation on slopes, we chose to vary the surface and ice properties in our sensitivity only within the longitudes of the Cthulhu region, between latitudes 40°N and 40°S. A complete description of the sensitivity of the CH4 and N2 cycles to these adjustable parameters can be found in Bertrand and Forget (2016); Bertrand et al. (2018) and Bertrand et al. (2019).
The baseline values were retained for the rest of the surface to maintain realistic pressure and CH4 cycles.
Table 1: Input parameters for the Pluto VTM and the values adopted in our sensitivity study. [^1] For the sensitivity study, the parameter values were only varied within the Cthulhu region, while elsewhere, they remained fixed at their baseline values. The transition between the albedo of the bare surface and that of a CH4-frosted surface follows the parameterization of Bertrand et al. (2020). The values listed in this table correspond to the maximum albedo that a frosted surface can reach, typically for deposits thicker than $\sim$ 1 mm. TI refers to the thermal inertia, the subscript d corresponds to the diurnal thermal inertia, i.e., the first centimeters of the surface, and the subscript s refers to the seasonal thermal inertia (depths deeper than $\sim$ 5 cm). The baseline values for the first five rows were adjusted to ensure a realistic representation of the seasonal N2 and CH4 cycles.
@@
For each set of parameters, the model was run for 100 Pluto years to equilibrate the subsurface temperatures on each subgrid slope, with slope condensation/sublimation disabled to avoid the formation or disappearance of thick glaciers due to non-equilibrated temperatures. We then ran the model for an additional ten years, allowing the condensation of volatiles on each subgrid slope. Only results from the final year are presented in this manuscript.
@@ 3.1 Frost mapping and statistics
Figure 4: CH4-integrated LEISA band depth maps from Schmitt et al. (2017) and Gabasova et al. (2021), overlaid with the LORRI albedo map (Hofgartner et al., 2023). The white patches in the Cthulhu region, which are not colored, are considered to be CH4 frost (see the main text for an explanation).
The geographic distribution of frost identified in this study, based on the methodology described in Sect. 2.1 is presented in Fig. 4. A comparison between the LEISA CH4 and N2 band depth maps at this location indicates that these frosts correspond to N2 ice deposits in Sputnik Planitia (northeastern part of Fig. 4) and Elliot crater (located around 12°N, 139°E), and to CH4 frost elsewhere. CH4 frost is detected poleward of 10°N, but its identification at lower latitudes is more difficult, as the bright spots observed with LORRI are not always associated with a distinct signal in the CH4 integrated LEISA band depth. This bias arises from the resolution difference between the LORRI observations ($\sim$ 300 m per pixel, Weaver et al., 2008; Hofgartner et al., 2023) and the LEISA CH4 band depth maps ($\sim$ 2.7 km at best, Gabasova et al., 2021).
Nonetheless, several lines of evidence support the interpretation that these bright deposits are indeed CH4 frost. First, in the largest bright patches, a few pixels exhibit high CH4 integrated band depth values (up to 0.4). In such cases, the correlation between bright deposits and CH4 frost is robust, suggesting that the absence of a correlation elsewhere is primarily due to resolution differences between the instruments. Second, Earle et al. (2018b) identified CH4 ice deposits on slopes between 15°S and 15°N using MVIC data, which offers higher spatial resolution ($\sim$ 1 km Earle et al., 2022). Third, analysis of RGB images at these locations reveals that some bright, high-albedo slopes (Supplementary Fig. 6 of Bertrand et al., 2020) are adjacent to similarly oriented, red, volatile-free slopes (Earle et al., 2022), indicating that the bright signatures are not solely due to illumination effects, but also to the presence of surface ice. We therefore conclude that the bright deposits observed in Fig. 4 within the Cthulhu region are CH4 frost. The LEISA phase-index map (see Supplementary Fig. 7b of Bertrand et al., 2020) suggests that these deposits are relatively pure CH4. Earle et al. (2022) further suggested that pure CH4 ice is observed on crater rims at low latitudes, while CH4 diluted with N2 may occur on the north-facing crater walls at lower elevations.
@@
Figure 5: Normalized count of the slope angle (A), the slope azimuth (B), and the projected slope $\mu$ (C) where CH4 frost is detected for all locations in Fig. 4 (gray histogram) and the Cthulhu region alone ($\pm$ 10° latitudes; blue histogram).
3.2 Simulation of the frost formation and evolution on the slopes
@@ 3.2.1 Baseline results
The seasonal evolution of CH4 and N2 frosts on flat surfaces, and on south and north-facing slopes, using the baseline model parameters described in Table 1, is shown in Fig. 6. These results highlight the significant influence of slope orientation on the condensation and sublimation of N2 and CH4. At high latitudes (Figs. 6a and 6b), the amount of N2 frost forming on 30° north-facing slopes is about five times higher (three times for CH4) than on flat terrain. Furthermore, our model predicts perennial accumulation on 30° south-facing slopes, whereas only seasonal frost is predicted on flat surfaces. At 20°N and at the equator (Figs. 6c–6h), N2 and CH4 frosts form almost exclusively on slopes, with only negligible amounts on flat terrain. In the southern hemisphere (10°S; Figs. 6g and 6h), frost is again predicted on south-facing slopes and flat surfaces during southern winter, but not during northern winter, while our model also predicts frost formation on north-facing slopes.
@@
These patterns can be explained by the differential solar illumination received by surfaces with different orientations. As shown by Bertrand et al. (2020, see their supplementary Figs. 8 and 9), 30° north-facing slopes in the northern hemisphere receive about 35% less insolation than flat terrain during winter and spring (and conversely for 30° south-facing slopes). Volatiles should therefore accumulate preferentially on these cold traps relative to warmer surfaces. The seasonal evolution of frost deposits also follows the illumination cycle: frost forms in late autumn and winter, when slopes are in shadow, and sublimates in spring and summer, when they are exposed to sunlight. An important asymmetry is observed, however: Frost never completely disappears from south-facing slopes, unlike north-facing slopes (although the latter is highly dependent on the assumed albedo; see Sect. 3.2.2). In contrast, Bertrand et al. (2020) suggested, based on illumination computations, that frost on south-facing slopes should disappear after southern spring, when these slopes warm. The main difference between our model and theirs is the inclusion of the seasonal CH4 cycle, which strongly affects the sublimation rate of CH4 frost. As shown by Bertrand and Forget (2016, see their Fig. 2), the maximum atmospheric CH4 mixing ratio during southern summer is five times higher than during northern summer, because the atmosphere is then ten times thinner and provides less dilution of methane. As expressed in Eq. 7, the sublimation rate of CH4 frost is highly sensitive to the CH4 mixing ratio. This asymmetry results in a sublimation flux up to 3-6 times lower during the southern summer than in the northern summer, thereby favoring the formation of perennial CH4 frost deposits on south-facing slopes. There is also a seasonal asymmetry in the atmospheric -and methane- density $\rho$, with larger values during the northern spring compared to the southern spring. This induces stronger sublimation fluxes in the northern spring on warm, north-facing slope, and, conversely, enhanced CH4 condensation and accumulation on cold, south-facing slopes at the same periods, again favoring the preservation of frost and the formation of a perennial deposit. Once such a perennial deposit forms, its high albedo reduces the absorbed energy and the annual mean surface temperature, which in turn promotes further frost accumulation, establishing a positive feedback. Finally, although the Pluto orbit is highly eccentric, surface temperatures show no discernible asymmetry, suggesting that present-day CH4 frost dynamics are governed mainly by atmospheric variations of pressure, density, and methane mixing ratio rather than by direct thermal effects. Indeed, as shown by Bertrand et al. (2020, see their Fig. S8), north and south-facing slopes receive nearly the same amount of solar insolation during spring and summer, when illuminated, and thus reach comparable temperatures. The reason is that the equinox and perihelion on Pluto occur close in time, which produces polar nights of equal length and summers of similar duration and intensity in both hemispheres (Earle et al., 2017). During other epochs, the timing of equinox relative to perihelion differs (Dobrovolskis et al., 1997). As a result, the high orbital eccentricity of Pluto produces substantial variations in seasonal illumination (Earle et al., 2017) which may more importantly impact the formation and sublimation of N2 and CH4 deposits .and mitigate the atmospheric effects described here (Bertrand et al., 2018, 2019).
@@
Figure 6: Seasonal evolution of frost on 30° south-facing slopes (blue), flat surfaces (red), and 30° north-facing slopes (black) on Pluto, shown for latitudes 40°N (A, B), 20°N (C, D), 0°N (E, F), and 10°S (G, H). For each latitude, N2 frost thickness is shown in the left panels, and CH4 frost thickness in the right panels. The dashed gray line marks the date of the New Horizons flyby. For cases with perennial accumulations, thickness values have been rescaled for clarity. The frost thicknesses shown here are zonal averages of the model outputs between 40°E and 150°E, corresponding to the Cthulhu region.
@@ 3.2.2 Model sensitivity
Figure 7 illustrates the sensitivity of our model results to the various parameters described in Table 1. As expected, increasing the ice albedo enhances the stability of frost (either N2 or CH4) during spring and summer (by about 20 Earth years for the case of N2, Fig. 7a, and $\sim$7 years for CH4, Fig. 7b). The stability of CH4 frost is highly sensitive to its albedo (c.f. the albedo feedback effect described in Bertrand et al., 2020): increasing the ice albedo from 0.71 to 0.73 can lead to the formation of perennial ice deposits. Increasing the emissivity of the ice also enhances frost stability with a similar magnitude (Figs. 7c and 7d), as it improves ice cooling and thus reduces the sublimation rate. Frost deposits are also sensitive to the thermal conductivity of the surface and underlying bedrock (here defined at depths greater than 5 cm), and thereby to thermal inertia. High thermal inertia surfaces are more conductive and release the heat stored during daytime/summer into the cold seasons, which reduces the amount of ice that can condense and hence decreases its stability. As highlighted by Eq. 7 and Fig. 7g, the condensation/sublimation rate of CH4 ice is strongly controlled by wind speed: strong winds promote deposition but also accelerate sublimation, while weak winds inhibit both condensation and sublimation. In the latter case, however, given the thinness of the deposits, frost disappears faster than under high-wind conditions. Finally, slope angle also affects frost stability: steeper slopes receive lower mean annual insolation, leading to lower surface temperatures and thus favoring the formation of thicker and more persistent frost deposits (Fig. 7h). Additional sensitivity tests (not shown) indicate that bare surface albedo and emissivity affect the amount of heat absorbed by the ground during the day and its cooling at night, respectively. These processes influence the diurnal formation of frost and potentially its seasonal persistence. The amount of CH4 in the atmosphere, and hence its transport, also affects the stability of frost deposits, as previously explained (see last paragraph).
@@
Our sensitivity experiments show that the timing of frost formation and disappearance is highly sensitive to the choice of key parameters, notably albedo, emissivity, diurnal thermal inertia, and regolith thermal inertia. These parameters have also been identified as critical for the evolution of volatile cycles across Pluto (e.g., Bertrand and Forget, 2016; Bertrand et al., 2018, 2019; Young et al., 2021). On Mars, the seasonality of CO2 and H2O frosts has been used to constrain the vertical stratification of the subsurface and its properties (notably the thermal inertia; Vincendon et al. 2010b, a). Such an analysis is not possible here because of the strong multi-parameter sensitivity of our problem and the fact that we only have a single snapshot of frost dynamics, which leaves the system poorly constrained. Interestingly, Fig. 7 shows that the sublimation rate of CH4 ice deposits is more sensitive to their albedo than to emissivity. In our simulations, at typical CH4 frost temperatures, the infrared radiative cooling flux of the surface is two to three times weaker than the peak daytime insolation, explaining why albedo is the main driver of the sublimation rate.
@@
Figure 7: Sensitivity of the N2 and CH4 frost thicknesses to the model inputs: A) N2 ice albedo; B) CH4 ice albedo; C) N2 ice emissivity; D) CH4 ice emissivity; E and F) thick lines: surface thermal inertia (first 5 cm), dashed lines: bedrock (below 5 cm) thermal inertia G) wind speed H) slope angles. All of the frost thicknesses are for a 30° north-facing slope, located at 20°N.